# Reducing Spatial Discretization Error on Coarse CFD Simulations Using an OpenFOAM-Embedded Deep Learning Framework

### v2

@@ 4.2 Deep learning architecture
The architecture of the modified inverse operator $\widetilde{\mathcal{I}}$ comprises seven different types of blocks. We should mention that we treat each velocity component independently, so we take three inputs at time $t$: each component of the mean velocity and the mean pressure.
@@
where $U$ and $D$ are the upwind and the downwind cell to $P$, respectively. The variable $r_{F}$ measures the ratio between the upstream and downstream gradients from face $F$. In this case, (18) assumes a 1D stream and a mesh with cells of the same size. Total Variation Diminishing schemes inspire the computation of this variable [44], commonly used due to their outstanding performance. The most important advantage of these schemes is that they preserve monotonicity, ensuring a stable and non-oscillatory solution. This feature is reached by monitoring the $r$ values, permitting the differencing scheme to act if the range of r is between $[0,2]$ and changing to the Upwind scheme otherwise. This justifies our second type of blocks, in which we clip the values of $r$ to this acting range. The third block is the normalization one. Normalization of the neural network’s inputs is a standard process that offers several benefits, like faster convergence or enhanced generalization. Moreover, it is also necessary in this case because the orders of magnitude of pressure and velocities are different. We normalize the velocities and the pressure in the following way:
$\bar{u}^{i}_{norm}$ $:=\frac{\bar{u}^{i}}{max(\bar{u}^{i})},$ (19)
@@
The next block is patch extraction, which consists of going through every internal face of the domain and extracting the values of the neighboring cells after fixing the desired dimension of the stencil. We established a stencil size of $4\times 3$ for vertical faces and $3\times 4$ for horizontal faces. Figure 5 depicts a graphical description of the stencils employed.
@@
Figure 6: Architecture’s sketch.
The neural networks $\mathcal{N}^{x}_{e}$ and $\mathcal{N}^{y}_{e}$ are composed of 4 hidden layers of size 53, 49, 45, and 41, while the network $\mathcal{N}_{G}$ has two hidden layers of size 31 and 32. Altogether, the trainable parameters of the models add up to 22,036. In the case of the activation function, we employ the hyperbolic tangent to all the present layers.
@@ 4.3.1 Interpolation accuracy
Since we are enhancing the numerical accuracy of the convective process using some modified interpolation weights $\widetilde{\textbf{w}}=(\widetilde{w}^{x},\widetilde{w}^{y})$ generated by our neural network, we need to guarantee that the interpolation is at least first-order accurate like it is done for typical polynomial interpolation. We ensure the accuracy required by imposing the following restrictions:
@@
This restriction is carried out following the technique described in [21, 22], which consists of applying an affine transformation $\widetilde{\textbf{w}}=\textbf{A}\textbf{x}+\textbf{b}$ to the unconstrained output x of the neural network. Starting from the linear system of equations $\textbf{C}\alpha=\delta$ constructed from the standard formula for generating the finite difference coefficients for a $N$-point stencil, we define the reduced constant matrix $\textbf{C}_{r}$ as the first $N-1$ rows of C. Thus, the matrix A is the nullspace of the reduced constraint matrix, and the vector b is an arbitrary valid set of interpolation weights. In this case, it equals the standard-finite difference solution $\alpha$. Thus, if we need $N$ interpolation weights, our neural network will generate $N-1$, and we obtain the remaining one using this affine transformation. We also explore other options, such as producing the $N$ unconstrained coefficients by the neural network and dividing each one by their sum. However, this alternative induced instabilities in the training process and produced initial coefficients further away from the optimum in the solution space.
@@ 4.3.2 Bounding neural network prediction on boundary patches
We realized the neural network produced unphysical predictions for some boundary patches, especially around the column. This phenomenon is related to the near-wall treatment applied to the turbulent quantities and the size of the cells of the first wall layer. Consequently, the wall function damps the solution on these cells, producing accurate results for short-term simulations even if the neural network prediction is completely wrong. However, this small error accumulates and arises during mid-term simulations, producing a wrong generalization. We decided to implement a 1D limiter to bound the prediction of our neural network in these types of patches, inspired by the limiters of TVD schemes. Thus, considering a face $F$ surrounded by its owner cell $o$ and its neighbor cell $n$, we define the lower $b_{-}$ and upper $b_{+}$ bounds of our limiting criteria, such that:
@@
where $\lambda$ is a constant for extending or narrowing the range of action of the limiter. We will check the mean face velocities predicted in the boundary patches by our neural network so that if they are out of the bounds described in (22), we will change to the Upwind scheme. Hence, we obtain our bounded mean face velocity $\bar{u}^{*}_{F}$ as:
@@
We apply this constraint in the prediction stage once the model is trained in an unconstrained way. Otherwise, the generalization in the surroundings of the column deteriorates. A correct selection of $\lambda$ is crucial since, in case of underestimating or overestimating this parameter, we will not be correcting enough the solution or spoiling the neural network’s prediction, respectively. We set $\lambda=0.3$ experimentally using the training data once the neural network is completely trained.
@@ 4.4 Training process
We built the dataset for training the model from a CFD simulation of the domain described in Section 2 simulated for 20 s. However, during the first 4 s, the flow shows a transient and undesirable behavior until the wake starts to be formed. For this reason, we decided to discard the first 100 timesteps, resulting in 400 training snapshots. We created the input-output entries of the dataset following the details described in  A, after projecting the data to the coarse dimension and after fixing the number of timesteps accumulated $T$, according to the multi-step loss function (15). Due to the center-to-face approach, we generate several samples equal to the number of internal faces from a snapshot of the complete domain. The number of epochs and training samples depends directly on the number of timesteps accumulated, which will be defined for each case in the results section. Algorithm 1 formally describes the multi-step training process.
We employed a batch size of 30 and a stochastic gradient descent optimizer, Adam [45], with an initial learning rate of $\eta^{0}=0.01$ and the rest of the parameters as default. Moreover, we used a callback to reduce the learning rate on the fly if the validation loss did not decrease after a certain number $Z$ of epochs.
@@ 4.5.1 Communication between programs
We work with two different pieces of software that must communicate during prediction and training. Moreover, they are written in different programming languages, making communication more complex. The critical step becomes during training since this process requires bidirectional communication, as depicted in Figure 7. The training process also requires fast communication because the dataflow path is repeated hundreds of thousands of times while optimizing the network parameters.
@@
It is worth mentioning that TensorFlow divides the entire dataset into mini-batches of data to exploit GPU parallelization. In our case, we ran both the TensorFlow and the OpenFOAM parts in CPU, but we parallelized the execution of the simulations that compose each batch using the OpenMP [52] package. That is why our batch size is not a power of two as usual since we are limited to the number of CPU threads.
@@ 4.5.2 Differentiation of OpenFOAM
As seen in Section 4.1, we can describe our method as a composition of the functions $\mathcal{I}_{c}$ and $\mathcal{F}_{c}$, while the operator $\mathcal{L}$ applied to this composition returns the loss, as described in Figure 8. To obtain the optimum parameters of the neural network, we need to compute the derivative of the loss w.r.t. the parameters of the network $\theta$, which in our case is defined as:
@@
We fill this gap using a discrete adjoint version of OpenFOAM developed by M. Towara [55]. This method applies AD by operator overloading to the program’s core instead of solving the adjoint equations as a continuous adjoint software. By evaluating the chain rule on a per-operation level through the whole solver execution, it calculates the derivative we need. Additionally, the code allows to efficiently evaluate Jacobian-vector (tangent/forward AD) and vector-Jacobian (reverse/adjoint AD) products, leading to an economical and adequate solution for our problem. Note the strong relation in concepts between adjoint AD and backpropagation in ML [24].
@@ 5.1 Sensitivity study to the reduction of mesh resolution
$F_{r}=\frac{m_{f}^{x}}{m_{c}^{x}}=\frac{m_{f}^{y}}{m_{c}^{y}}$ (25)
where $m$ is the number of cells. As the reduction between meshes is isotropic, the reduction factor is the same in the directions of $x$ and $y$. We conducted the sensitivity study varying the resolution factor and observing the error by the multi-step mean absolute percentage error $\Psi$ under the same training conditions. The training with each mesh predicts just one timestep, so we fix the number of accumulated time steps $T$ to 1, obtaining 399 training snapshots. We establish the number of epochs or training iterations to 100, with a callback to reduce the learning rate by a factor of 0.5 if the loss does not decrease after six epochs.
Figure 9 (a) and Figure 9 (b) gather the evolution of the error of each component of the velocity of the training data once the training process is completed. Regarding the effect of the resolution factor, we observe a lower degradation in the prediction of the Deep-Convection model compared with the baseline solver for the x-component of the velocity. A similar degradation is observed for the y-component. For a resolution factor of ×$\times 4$, the reduction in the loss for the $\bar{u}_{x}$ is 45% and for $\bar{u}_{y}$ 37%, while in the case of ×$\times 32$, the x-component exhibits a reduction of 70% and the y-component 39%. These percentages may be controlled by varying the weighting factors of (16).
@@
In Figure 9 (c), we show the computational time required to execute one epoch of the training $t_{epoch}$. We observe that this computation time decreases drastically with the increase of the resolution factor. This makes the training for a reduction of 2x unfeasible due to its computational cost and the feasibility of the 4x case dependent on the available resources. For this reason, to test the model in the worst conditions, we decided to continue the experiments with the mesh with an 8x reduction factor.
One advantageous feature of the model is the fast convergence shown during training. As depicted in Figure 10, the Deep-Convection model outperforms the baseline solver after a couple of epochs for both velocity components. Moreover, it reaches 81% and 71% of the total loss reduction for x and y components within the first ten iterations. In contrast to standard deep learning models, which usually need more than 1000 epochs to complete the training, our model achieves convergence in around 70 epochs. This feature is vital when training with more than one accumulated timestep because it permits a reduction in the duration of the training. Regarding Figure 10, we also observe the good behavior of the minimization algorithm, producing a smooth reduction in the loss. Thus, we can state the correct reconstruction of the missed derivative, described in Section 4.5.2, and the successful coupling between TensorFlow and OpenFOAM.
@@ 5.2 Performance on mid-term simulation
In the previous section, we observed promising results from the model predicting one single timestep. However, it is necessary to evaluate the mid-term performance of the model since stability and correct generalization of the training data are crucial for obtaining a valuable method. For this purpose, we trained the model from 1 to 6 accumulated timesteps until convergence and from 7 to 13 with a fixed number of epochs. We also employed the callback mentioned in the previous section to reduce the learning rate on the fly.
@@
The experiment consists of a simulation of the training data from the first to the last. Thus, the baseline solver and the Deep-Convection start from the same initial conditions at 4 s and evolve until reaching 20 s, going along 399 timesteps. It is evident that the longer the simulation, the higher the effect of the coarse mesh in the error since the error accumulates from one timestep to another. Furthermore, the relative error in the y-component of the velocity will be higher than in the x-component due to the manifestation of the von Karman effect as an oscillation along the y-axis. Figure 11 gathers the evolution of all the variables involved in the simulation.
@@
Figures 11(a) and 11(b) show that the Deep-Convection model reduces the error more than 50% both in the x- and y-velocity components until $t=12$ s. The reduction is even higher for the following timesteps, reaching a reduction above 75% of both velocity components at $t=20$ s. This improvement in the performance of the DC model after $t=12$ s is due to the full development of the flow at this point, being the under-developed flow harder to enhance. The relative error in the y-component of the velocity is almost five times higher than in the x-component, as anticipated, and the increase of the error with time is much more controlled with the DC model than with the baseline one. At $t=20$ s, the velocity results are completely different, being the ones produced with the DC model more similar to the ground truth, as depicted in Figure 12.
@@
Figure 12: Comparison of each component of the velocity field at $t=20$ s between the ground true, the Deep-Convection model, and the baseline solver.
@@
Figures 11(d) and 11(e) show a similar behaviour of the turbulent quantities $k$ and $\omega$, reducing the error from 70% to 20% and from 50% to 25% at $t=20$ s, respectively. Like the velocity, the error reduction increases after $t=12$ s, when the flow becomes a fully developed flow. Hence, the turbulent variables modeling improves indirectly by enhancing the velocity with fine-scale effects. Figure 11(c) depicts that the pressure also reduces its error, but the reduction is far less significant. Before $t=12$ s, the pressure error of the DC model is practically the same as the baseline one, increasing after this point until around 20%.
Although the reduction in pressure is minor, we observe an evident greater similitude of the field produced by the DC model with the ground truth than one produced by the baseline model, as depicted in Figure 13. In our case, the intention is to accurately reconstruct the weak’s velocity; consequently, the approach works correctly. However, if we aim to reduce the error of the pressure field, for example, to approximate the drag coefficient at the column in the coarse mesh to the finer one, we should add the pressure contribution to the loss function. Alternatively, an extra neural network could be added to correct the pressure field.
@@
Figure 13: Comparison of pressure field at $t=20$ s between the ground true, the Deep-Convection model, and the baseline solver.
@@
Model Computation time per timestep [s]
@@
The difference in computation time between the baseline model using the coarse mesh and the DC model is due to the deep-learning framework. This simulation time can be drastically reduced by employing the graph TensorFlow mode instead of the eagerly one we employed. In our case, the eagerly mode is justified in facilitating the development of the communication code between Python and C++. Despite this penalty, the DC model is 60 times faster than the fine baseline model. Thus, the results Figure 11 shows take about 4 minutes, with a 25% error compared with the high-precision simulation, which takes 4 hours in a single core. The time for training the model was 38.11 h, although the code can also be further optimized. Based on experiments in simpler cases, we estimate a reduction 35% in training time could be possible by code optimization, added to the reduction obtained employing the graph TensorFlow mode.

### v3

@@ 4.2 Deep learning architecture
The architecture of the modified inverse operator $\widetilde{\mathcal{I}}$ comprises seven different types of operations, which we named as blocks, and we depicted their connectivities in Figure 6. We treat each velocity component independently, so we take three inputs at time $t$: each component of the mean velocity and the mean pressure.
Figure 6: Architecture’s sketch.
@@
where $U$ and $D$ are the upwind and the downwind cell to $P$, respectively. The variable $r_{F}$ measures the ratio between the upstream and downstream gradients from face $F$. In this case, (18) assumes a 1D stream and a mesh with cells of the same size. Total Variation Diminishing schemes inspire the computation of this variable [48], commonly used due to their outstanding performance. The most important advantage of these schemes is that they preserve monotonicity, ensuring a stable and non-oscillatory solution. This feature is reached by monitoring the $r$ values, permitting the differencing scheme to act if the range of r is between $[0,2]$ and changing to the Upwind scheme otherwise. The $r$ variable can be interpreted as a second derivative or as a measure of the acceleration of the velocity field at $F$. Values below zero or higher than two mean a sharp change in the acceleration, being the face velocity totally influenced by the upwind or the downwind cell, respectively. For values between zero and two, our network should discover how to combine the neighboring cell center values. This interpretation justifies our second type of blocks, that clip the values of $r$ to the acting range $[0,2]$. The third block, following Figure 6, is the normalization one. Normalization of the neural network’s inputs is a standard process that offers several benefits, like faster convergence or enhanced generalization. Moreover, it is also necessary in this case because the orders of magnitude of pressure and velocities are different. We normalize the velocities and the pressure in the following way:
$\bar{u}^{i}_{norm}$ $:=\frac{\bar{u}^{i}}{max(\bar{u}^{i})}\qquad\text{with $i\in\{x,y\}$},$ (19)
@@
The next block, (see Figure 6), is patch extraction, which consists of going through every internal face of the domain and extracting the values of the neighboring cells after fixing the desired dimension of the stencil. We establish a stencil size of $4\times 3$ for vertical faces and $3\times 4$ for horizontal faces. Figure 7 depicts a graphical description of the stencils employed.
@@
Afterward, the two groups of patches are sent, as Figure 6 depicts, respectively to two neural networks $\mathcal{N}^{x}_{e}$ and $\mathcal{N}^{y}_{e}$ of the same shape, which act as encoders compressing the patch information ($r_{x},r_{y},\bar{u}^{x},\bar{u}^{y},\bar{p}$) to a latent space. We gather these networks’ output and insert them into a third neural network $\mathcal{N}_{G}$, which generates the modified interpolation weights $\widetilde{\textbf{w}}$. Finally, in the last block, called $\textbf{u}_{F}$ computation, we obtain the mean face velocities based on the original mean velocities following (11). Figure 6 outlines the blocks defined and their connectivities.
The neural networks $\mathcal{N}^{x}_{e}$ and $\mathcal{N}^{y}_{e}$ are composed of four hidden layers of size 53, 49, 45, and 41, while the network $\mathcal{N}_{G}$ has two hidden layers of size 31 and 32. Altogether, the trainable parameters of the models add up to 22,036. In the case of the activation function, we employ the hyperbolic tangent to all the present layers. We initialize these neural networks following the standard Keras initialization, namely, the weights employ a uniform Xavier/Glorot initialization [49] and the bias are set to zero.
@@ 4.3.1 Interpolation accuracy
We enhance the numerical accuracy of the convective process using some modified interpolation weights $\widetilde{\textbf{w}}=(\widetilde{w}^{x},\widetilde{w}^{y})$ generated by our neural network; thus, we need to guarantee that the interpolation is at least first-order accurate like it is done for typical polynomial interpolation. We ensure the accuracy required by imposing the following restrictions:
@@
We impose this restriction following the technique described in [21, 22], which applies an affine transformation $\widetilde{\textbf{w}}=\textbf{A}\textbf{x}+\textbf{b}$ to the unconstrained output x of the neural network. Starting from the linear system of equations $\textbf{C}\alpha=\delta$ constructed from the standard formula for generating the finite difference coefficients for a $N$-point stencil, we define the reduced constant matrix $\textbf{C}_{r}$ as the first $N-1$ rows of C. Thus, the matrix A is the nullspace of the reduced constraint matrix, and the vector b is an arbitrary valid set of interpolation weights. In this case, it equals the standard-finite difference solution $\alpha$. Thus, if we need $N$ interpolation weights, our neural network will generate $N-1$, and we obtain the remaining one using this affine transformation. We also explore other options, such as producing the $N$ unconstrained coefficients by the neural network and dividing each one by their sum. However, this alternative induced instabilities in the training process and produced initial coefficients further away from the optimum in the solution space.
@@ 4.3.2 Bounding neural network prediction on boundary patches
In a learning model predicting temporal data, we distinguish three different regions: short-, mid-, and long-term simulations. The first region encompasses those simulations inside the training distribution, which the model should perform successfully. Mid-term simulations are those outside the training distribution but close to the seen data, so a well-designed model must predict them correctly. Long-term simulations are related to predictions far away from the training distribution, where a learning model just based on data is incapable of being accurate [17] due to the weak extrapolating capabilities of neural networks.
@@
The the neural network can produce unphysical predictions for some boundary patches, especially around the column. This phenomenon is related to the near-wall treatment applied to the turbulent quantities and the size of the cells of the first wall layer. Consequently, the wall function damps the solution on these cells, producing accurate results for short-term simulations even if the neural network prediction is completely wrong. However, this small error accumulates and arises during mid-term simulations, producing a wrong generalization. Thus, we implement a 1D limiter to bound the prediction of our neural network in these types of patches, inspired by the limiters of TVD schemes. Thus, considering a face $F$ surrounded by its owner cell $o$ and its neighbor cell $n$, we define the lower $b_{-}$ and upper $b_{+}$ bounds of our limiting criteria, such that:
@@
where $\lambda$ is a constant for extending or narrowing the range of action of the limiter. We check the mean face velocities predicted in the boundary patches by our neural network so that if they are out of the bounds in (22), we change to the Upwind scheme. Hence, we obtain our bounded mean face velocity $\bar{u}^{*}_{F}$ as:
@@
We apply this constraint in the prediction stage once the model is trained in an unconstrained way. Otherwise, the generalization in the surroundings of the column deteriorates. A correct selection of $\lambda$ is crucial since, in case of underestimating or overestimating this parameter, we will not be correcting enough the solution or spoiling the neural network’s prediction, respectively. We set $\lambda=0.3$ experimentally using the training data once the neural network is completely trained. Besides, we employ the same mesh description as OpenFOAM, so we can easily filter and detect boundary patches.
@@ 4.4 Training process
We built the dataset for training the model from a CFD simulation of the domain described in Section 2 simulated for 20 s. However, during the first 4 s, the flow shows a transient and undesirable behavior until the wake starts to be formed. For this reason, we discard the first 100 timesteps, resulting in 400 training snapshots. We create the input-output entries of the dataset following A, after projecting the data to the coarse dimension and after fixing the number of timesteps accumulated $T$, according to the multi-step loss function (15). Due to the center-to-face approach, we generate a number of samples equal to the number of internal faces from a snapshot of the complete domain. The number of epochs and training samples depends directly on the number of timesteps accumulated, which will be defined for each case in the results section. Algorithm 1 formally describes the multi-step training process.
We employ a batch size of 30 and a stochastic gradient descent optimizer, Adam [50], with an initial learning rate of $\eta^{0}=0.01$ and the rest of the parameters as default. Moreover, we use a callback to reduce the learning rate on the fly if the training loss did not decrease after a certain number $Z$ of epochs. We monitor the training convergence by a stopping criterion that checks if the loss function plateaus during a certain number of iterations (around 20), once we reach the lowest learning rate we set in the callback.
@@ 4.5.1 Communication between programs
We work with two different pieces of software that must communicate during prediction and training. Moreover, they are written in different programming languages, making communication more complex. The critical step becomes during training since this process requires bidirectional communication, as Figure 8 depicts. The training process also requires fast communication because the dataflow path is repeated hundreds of thousands of times while optimizing the network parameters.
@@
It is worth mentioning that TensorFlow divides the entire dataset into mini-batches of data to exploit GPU parallelization. In our case, we ran both the TensorFlow and the OpenFOAM parts in CPU, but we parallelized the execution of the simulations that compose each batch using the OpenMP [57] package. That is why our batch size is not a power of two as usual since we are limited to the number of CPU threads. In the next version of the code, we plan to make the deep learning part fully GPU-compatible, so that the execution of the code becomes a sequential GPU-CPU-GPU procedure.
@@ 4.5.2 Differentiation of OpenFOAM
In Section 4.1, we describe our method as a composition of the functions $\mathcal{I}_{c}$ and $\mathcal{F}_{c}$, while the operator $\mathcal{L}$ applied to this composition returns the loss, as Figure 9 describes. We obtain the optimum parameters of the neural network by computing the derivative of the loss w.r.t. the parameters of the network $\theta$, which in our case is defined as:
@@
We fill this gap using a discrete adjoint version of OpenFOAM developed by Towara [60]. This method applies AD by operator overloading to the program’s core instead of solving the adjoint equations as a continuous adjoint software. By evaluating the chain rule on a per-operation level through the whole solver execution, it calculates the derivative we need. Additionally, the code allows to efficiently evaluate Jacobian-vector (tangent/forward AD) and vector-Jacobian (reverse/adjoint AD) products, leading to an economical and adequate solution for our problem. Note the strong relation in concepts between adjoint AD and backpropagation in ML [24].
@@ 5.1 Sensitivity study to the reduction of mesh resolution
$F_{r}=\frac{m_{f}^{x}}{m_{c}^{x}}=\frac{m_{f}^{y}}{m_{c}^{y}},$ (25)
where $m$ is the number of cells. As the reduction between meshes is isotropic, the reduction factor is the same in the directions of $x$ and $y$. We conducted the sensitivity study varying the resolution factor and observing the error by the multi-step mean absolute percentage error $\Psi$ under the same training conditions. The training with each mesh predicts just one timestep, so we fix the number of accumulated time steps $T$ to 1, obtaining 399 training snapshots. We establish the number of epochs or training iterations to 100, with a callback to reduce the learning rate by a factor of 0.5 if the loss does not decrease after six epochs. The obtained weighting factors (described in (16)) were almost constant for each of the coarse discretizations analysed. For this reason and for carrying out a fair comparison, we employed the same weighting factors in this sensitivity analysis.
@@
Figures 10(a) and 10(b) gather the resulting value of the loss function of each component of the velocity once the training process is completed (last epoch value). Regarding the effect of the resolution factor, we observe a lower degradation in the prediction of the Deep-Convection model compared with the baseline solver for the x-component of the velocity. A similar degradation is observed for the y-component. For a resolution factor of ×$\times 4$, the reduction in the loss for the $\bar{u}_{x}$ is 45% and for $\bar{u}_{y}$ 37%, while in the case of ×$\times 32$, the x-component exhibits a reduction of 70% and the y-component 39%. These percentages may be controlled by varying the weighting factors of (16).
Figure 10(c) shows the computational time required to execute one epoch of the training $t_{epoch}$. We observe that this computation time decreases drastically with the increase of the resolution factor. This makes the training for a reduction of x2 unfeasible due to its computational cost and the feasibility of the x4 case dependent on the available resources. For this reason, to test the model in the worst conditions, we decided to continue the experiments with the mesh with an x8 reduction factor.
One advantageous feature of the model is the fast convergence shown during training. Figure 11 shows how the Deep-Convection model outperforms the baseline solver after a couple of epochs for both velocity components. Moreover, it reaches 81% and 71% of the total loss reduction for x and y components within the first ten iterations. In contrast to previous “domain-based” models presented in Section 1, which usually need over 1000 epochs to complete the training, our “patch-based” model achieves convergence in around 70 epochs. This feature is vital when training with more than one accumulated timestep because it permits a reduction in the training duration. The rest of the resolution cases show a similar training curve, with almost identical speed of convergence. Regarding Figure 11, we also observe the good behavior of the minimization algorithm, producing a smooth reduction in the loss. Thus, we can state the correct reconstruction of the missed derivative, described in Section 4.5.2, and the successful coupling between TensorFlow and OpenFOAM.
@@ 5.2 Performance on mid-term simulation
The promising results of the previous section show that the model can predict a single timestep effectively. However, it is necessary to evaluate the mid-term performance of the model since stability and a correct generalization are crucial for obtaining a valuable method. Thus, we trained the model from 1 to 6 accumulated timesteps until convergence and from 7 to 13 with a fixed number of epochs. Generalization properties can be improved by training with a large value of T. Unfortunately, as T increases, the training costs also increase, so it is necessary to find a balance between generalization and cost. In addition, the increase in T must be progressive. Otherwise, the training will diverge. In the first stages of the training with $1\leq T\leq 6$, we can execute as many epochs as we want since the cost is affordable, so we extend the execution until convergence. For $T\geq 6$ the execution costs increase. However, by executing a limited number of epochs (without reaching full convergence), the method retains adequate generalization properties and accuracy since the highest loss decrease occurs within the first training iterations (see Figure 11). We also employed the callback mentioned in the previous section to reduce the learning rate on the fly.
@@
The experiment simulates of the training data from the first to the last. Thus, the baseline solver and the Deep-Convection start from the same initial conditions at 4 s and evolve until reaching 20 s, going along 399 timesteps. It is evident that the longer the simulation, the higher the effect of the coarse mesh in the error since the error accumulates from one timestep to another. Furthermore, the relative error in the y-component of the velocity is higher than in the x-component due to the manifestation of the von Kármán effect as an oscillation along the y-axis. Figure 12 gathers the evolution of all the variables involved in the simulation.
@@
Figures 12(a) and 12(b) show that the Deep-Convection model reduces the error more than 50% both in the x- and y-velocity components until $t=12$ s. The reduction is even higher for the following timesteps, reaching a reduction above 75% of both velocity components at $t=20$ s. This improvement in the performance of the DC model after $t=12$ s is due to the full development of the flow at this point, being the under-developed flow harder to enhance. The relative error in the y-component of the velocity is almost five times higher than in the x-component, as anticipated, and the increase of the error with time is much more controlled with the DC model than with the baseline one. At $t=20$ s, the velocity results are completely different, being the ones produced with the DC model more similar to the ground truth, as depicted in Figure 13.
@@
Figure 13: Comparison of each component of the velocity field and error at $t=20$ s between the ground truth, the Deep-Convection model, and the baseline solver.
@@
Figures 12(d) and 12(e) show a similar behaviour of the turbulent quantities $k$ and $\omega$, reducing the error from 70% to 20% and from 50% to 25% at $t=20$ s, respectively. Like the velocity, the error reduction increases after $t=12$ s, when the flow becomes a fully developed flow. Hence, the turbulent variables modeling improves indirectly by enhancing the velocity with fine-scale effects. Figure 12(c) depicts that the pressure also reduces its error, but the reduction is far less significant. Before $t=12$ s, the pressure error of the DC model is practically the same as the baseline one, increasing after this point until around 20%.
Although the improvement in the pressure modeling is less significative than in the rest of the variables, we observe a greater similitude between the ground truth vs. the field produced by the DC model than vs. the field of the baseline model, as Figure 14 shows. Another model limitation comes from the computation of hydrodynamic coefficients. Due to the problem caused by the law of the wall function (see Section 4.3.2), the model predicts drag coefficients with no enhancement compared with the baseline model for timesteps beyond those seen during training. For this reason, our model is inadequate for computing the temporal evolution of drag or lift coefficients. A solution to this problem could be to add an extra neural network to correct the cells affected by the law of the wall function, switch to a LES simulation, or impose the hydrodynamic coefficients as hard constraints in the loss function. This topic will be studied in upcoming works.
@@
Figure 14: Comparison of pressure field at $t=20$ s between the ground truth, the Deep-Convection model, and the baseline solver.
@@
Model Computation time per coarse timestep [s]
@@
The difference in computation time between the baseline model using the coarse mesh and the DC model is due to the deep-learning framework. This time difference can be drastically reduced by employing the graph TensorFlow mode instead of the eager mode we employed, making it almost negligible. In our case, the eager mode is justified in facilitating the development of the communication code between Python and C++. Despite this penalty, the DC model is 60 times faster than the fine baseline model. Thus, the results Figure 12 shows take about 4 minutes, with a 25% error compared with the high-precision simulation, which takes 4 hours in a single core. The time for training the model (excluding the time for generating the training data) was 38.11 h, although the code can also be further optimized. Based on experiments in simpler cases, we estimate a reduction 35% in training time could be possible by code optimization, added to the reduction due to the graph TensorFlow mode. The time employed in converting the fine data to the coarse dimension and in constructing the dataset from the CFD simulation is negligible compared with the training time, being it in the order of seconds.